OpenStudy (anonymous):

Find y" in terms of x and y by implicit differentiation. (x^3)+(y^3)=4^3

OpenStudy (anonymous):

I think the first would start out as 3y^2(dy/dx) so 3x^2+3y^2(dy/dx)=0

OpenStudy (anonymous):

Dang, what am I doing wrong, the professor didn't really go over this rule in my class.

OpenStudy (anonymous):

ok thats right carry on sorryyyyyyyyyyyyyy

OpenStudy (anonymous):

so would the second derivative be 6x+6y(dy/dx)=0

OpenStudy (anonymous):

ya right

OpenStudy (anonymous):

Okay then after that how would you put it in terms of x and y? I am guessing to start you would put all the y's on one side like this maybe? 6x=6y(dy/dx)

OpenStudy (anonymous):

ya but 6x= -6y(dy/dx) => dy/dx = -x/y

OpenStudy (anonymous):

but remember this is the second derivative

OpenStudy (anonymous):

Now I am stuck, sorry.

OpenStudy (anonymous):

why where????

OpenStudy (anonymous):

I am stuck on simplifying it down to terms x and y

OpenStudy (anonymous):

ok ok see this

OpenStudy (anonymous):